Given a{n}=-(-1) u[n – 1] H[n]= a[n]+(6) a[n-1] find y[n] = x[n]*h[n].

note: if you get repeated roots, you're doing it wrong.
Answer #1

Given that, X(n) =-(-) u (n.) hin) = s(n) + ( 2 ) = u(n-1) Convolution : y(n) = x(n) *h(n) = [(-luc-in) + (scn) + ( ) (-1) 1-aucul uln) = ® -(44) un-) + (5) Cun-) considel - (- Jurn) & (Sun) - Considdi anti- a-b uln) - cyc.) + (%lum) + (-4)*- (4) Ju